# Correspondences of the Third Kind: Camera Pose Estimation from Object Reflection

### v1

@@ 3.5 Camera Pose Estimation from Two Images
Figure 3 depicts how the camera pose between two views can be recovered. To recover the normal maps ${\bf N}_{1}$ and ${\bf N}_{2}$, we use DeepShaRM which is a radiometry-based geometry estimation method [21] to each input image independently. This gives us the normal maps but only up to the GBR transformations. DeepShaRM also recovers a reflectance map for each view from the normal maps. We use these outputs for the reflection correspondence detection.
To obtain 3D correspondences in the normal maps, we train a deep network to extract pixel-wise, view-invariant features that we use to match locations across the normal maps. Figure 5 depicts this feature extraction network. We train the network by contrastive learning. We use pairs of synthetic normal maps of two views as training data. We feed both of them to the feature extraction network and obtain corresponding feature maps ${\bf F}_{1}$ and ${\bf F}_{2}$. Given ground truth correspondences $(u_{1}^{i},v_{1}^{i})\leftrightarrow(u_{2}^{i},v_{2}^{i})$, we impose the InfoNCE loss [17]
@@
The training with synthetic normal maps is, however, insufficient for 3D correspondence detection in the estimated normal maps as they are distorted by GBR transform. We overcome this with data augmentation. For each view in the training data, we randomly sample parameters of the GBR transformation and transform the input normal map according to Eq. 4. We use the transformed normal maps instead of the original ones so that the network can learn to extract features robust to the estimation errors caused by the bas-relief ambiguity.
At inference time, using the extracted feature maps, we detect correspondences by brute-force matching and filter them with the ratio test by Lowe [10]. For any location on the object in the first image, we look for the best and second best locations in the second image in terms of the cosine similarity of the features. If the ratio of the similarity of the second best location to one of the best location is lower than a threshold, we use the best location as the matched location. This gives us both pixel correspondences and 3D correspondences.
We also train another deep neural network reflection correspondences in the same way. Please see the supplementary for more details.
Figure 4: Once we recover the camera poses, we can iteratively refine them by alternating between estimation of the surface normals and reflectance maps, and the camera poses. We use DeepShaRM [21] to update the normals and reflectance maps. DeepShaRM also outputs a unified surface geometry estimate as a 3D grid of a singled distance function (SDF).
Once we have the three types of correspondences, we recover the relative rotation using our two-step algorithm (Sec. 3.4). For robust estimation, we perform the first step with RANSAC [5]. We estimate different candidates of the combined transformation $\mathrm{G_{21}}$ from randomly sampled minimal sets of pixel and 3D correspondences. We then select one of them according to the number of inlier correspondences, i.e., ones that are consistent with the estimated parameters. In the second step, for each possible decomposition of $\mathrm{G_{21}}$, we check the number of inlier reflection correspondences that are consistent with the decomposed parameters and select one according to it. Based on Sec. 4.1, we use four pixel and four 3D correspondences from the normal map pair as the minimal set for the estimation in the first step.
3.5.1 Translation Estimation
Once $\mathrm{R_{21}}$ is determined, by leveraging the pixel correspondences, we can also solve for the relative translation between the two views except for offsets regarding the two viewing directions [6]. That is, we estimate the translation vector $\mathbf{t_{21}}=(t_{x},t_{y},t_{z})$ using the constraint
where $r_{jk}$ is $(j,k)$ element of $\mathrm{R_{21}}$ and
$\tilde{u}^{i}=r_{11}u_{2}^{i}+r_{12}v_{2}^{i}-u_{1}^{i}\,,$ (23)
$\tilde{v}^{i}=r_{21}u_{2}^{i}+r_{22}v_{2}^{i}-v_{1}^{i}\,.$ (24)
Figure 5: We train a deep neural network to extract view-independent, per-pixel features for correspondence matching. Contrastive learning using InfoNCE loss [17] with data augmentation with the GBR transform enables feature extraction for distortion-robust correspondence matching.
@@ 3.6 Joint Shape and Camera Pose Recovery
As depicted in Fig. 4, we further improve the camera pose estimation accuracy by alternating between estimating camera poses using the surface normal and reflectance map estimates and updating them using the camera pose estimates. As a byproduct, we obtain an accurate object shape. For the update of the surface normal and reflectance map estimates, we use DeepShaRM [21]. Using the camera poses as inputs, DeepShaRM integrates the multi-view estimates of surface normals and output improved estimates along with a surface geometry estimate represented as a 3D signed distance function.
If we have more than two views, we can integrate the estimated relative camera pose for each view pair. We first initialize a camera pose estimate for each view $\mathrm{R_{k}}$ using some of the recovered relative poses. Using inlier correspondences, we then refine the pose estimates by minimizing the sum of the objective functions from Eq. 13 over each view pair.
@@ 4 Experimental Results
Figure 6: Empirical analysis of the number of correspondences required for optimization of the camera pose and the two GBR transforms of two images. See text for details.
@@
w/o Data Augm. 13.4 deg
w/o Joint 7.1 deg
Ours 6.1 deg
Table 1: Camera pose estimation accuracy on images in the nLMVS-Synth dataset [20]. Please see the text for details.
We focus our experiments on answering the following key questions. What are the actual number of correspondences required for resolving the bas-relief ambiguity? Can the novel feature extraction network establish correspondences robust to GBR distortion? To answer these questions, we validate the following points
  • The intuition about the number of required correspondences is consistent with results of empirical evaluation.
@@
  • The proposed components (e.g., the data augmentation) are essential for accurate reconstruction from a sparse set (2) of images.
4.1 The Number of Correspondences
(a) 3D Correspondences
(b) Reflection Correspondences
Figure 7: Examples of 3D and reflection correspondences detected by our method. Visual inspection shows that they match semantically correct surface and reflectance map points.
Figure 8: Surface normal and reflectance maps recovered from two synthetic images in the nLMVS-Synth dataset. The bas-relief ambiguity is successfully resolved and the recovered maps are consistent with ground truth.
Figure 9: Surface geometry recovered by the joint estimation.
We first conduct empirical evaluation of the number of correspondences required for resolving the bas-relief ambiguity for each view pair. We created 10000 sets of synthetic correspondences by randomly sampling surface normal orientations, relative rotation matrices, and parameters of the GBR transformation. Using this synthetic data, we checked how the numbers of correspondences used for the two-step algorithm (Sec. 3.4) affects the estimation accuracy. In this experiment, we optimized all the unknown parameters with all available correspondences using an off-the-shelf solver [18] as the two step algorithm cannot be applied to cases where reflection correspondences are not available. In the following results, we exclude cases where the objective function did not converge to zero as it is a symptom that the optimization algorithm did not converge correctly.
The blue line in Fig. 6 shows the relationship between the number of correspondences in normal maps (pixel and 3D correspondences) and rate of failure cases where the error (Frobenius norm) between estimated combined transformation matrix $\mathrm{G_{21}}$ and the ground truth is higher than 0.01. To simulate the first step of the algorithm, we did not use reflection correspondences in this experiment. The results show that four correspondences are required for the estimation of $\mathrm{G_{21}}$. The orange line in Fig. 6 shows the relationship between the number of reflection correspondences and rate of failure cases where the error between the estimated camera pose and the ground truth is higher than 0.1 degree. In this experiment, we used four pixel and four 3D correspondences in addition to the reflection correspondences. The results show that one reflection correspondence resolves the remaining ambiguity. Note that, the leftmost point in the orange line shows that, without reflection correspondences, we cannot decompose the combined transformation $\mathrm{G_{21}}$. These results are consistent with the intuition discussed in Sec. 3.4.
4.2 Camera Pose Estimation
We evaluate the accuracy of the joint camera pose and object shape estimation framework on synthetic images. We train the feature extraction network using synthetic normal and reflectance maps on the nLMVS-Synth dataset [20].
@@ 4.2 Camera Pose Estimation
For quantitative evaluation, we test our method on synthetic images from the test set of the nLMVS-Synth dataset [20] which provides ground truth data. We compare our camera pose estimation accuracy (the average pose estimation error) with those by its own ablated variants, “w/o Data Aug.” and “w/o Joint”. “w/o Data Aug.” uses the feature extraction networks trained without the data augmentation with GBR transforms. “w/o Joint” is our method without the joint iterative estimation which recovers camera poses from only the initial estimates of normal and reflectance maps.
@@
Table 1 shows the average camera pose estimation errors. The results show that the joint estimation improves the estimation accuracy and the data augmentation method is essential for joint estimation of the relative camera pose and the parameters of the bas-relief transformation.
Figure 7 shows inlier correspondences detected by our method. The results join points on the 3D surface and reflectance maps that visually appear to be correct correspondences, demonstrating the effectiveness of the feature extractor and the robust estimation method. Their accuracy is ultimately reflected in the accuracy of the resulting camera pose estimates.
Figure 8 shows how the initial estimates of surface normal and reflectance maps are improved by the joint estimation of geometry and camera poses. Even when the initial estimates are erroneous, our method successfully resolves the bas-relief ambiguity and recovers plausible surface normal and reflectance maps along with the relative camera pose. As shown in Figure 9, surface geometry is also successfully recovered.
Please also see the supplementary material for quantitative and qualitative evaluation on real-world images and comparison with existing methods for camera pose estimation and joint shape estimation [3, 16].
@@ 5 Conclusion
We introduced a new type of correspondences, reflection correspondences, that enables accurate and robust camera pose and joint shape estimation from the appearance alone of textureless, reflective objects. We believe this new type of correspondences not only expunges restricting requirements on image capture for camera pose estimation but also opens new use of object appearance in various applications.
Acknowledgement
KY was in part supported by JST JPMJSP2110 and JPMJCR20G7. VL was in part supported by the ANR-DFG-JST TOSAI project and the ERC grant ”explorer” (No. 101097259). KN was in part supported by JSPS 20H05951, 21H04893, JST JPMJCR20G7, and RIKEN GRP.
Supplementary Material
Appendix A Ambiguity in The Estimation with Pixel and 3D Correspondences
@@
$\min_{z_{2}^{i},t_{x},t_{y}}f_{\mathrm{IM}}^{\prime}=\min_{z_{2}^{i},t_{x},t_{y}}\sum_{i}f_{u}^{i}+f_{v}^{i}\,,$ (A.1)
where $z_{2}^{i}$ is the depth of a surface point in the second view for the $i$-th correspondence, $t_{x}$ and $t_{y}$ are x and y elements of the relative translation vector, respectively, $f_{u}^{i}$ and $f_{v}^{i}$ are reprojection errors along the x and y axes, respectively,
$f_{u}^{i}=(r_{11}u_{2}^{i}+r_{12}v_{2}^{i}+r_{13}z_{2}^{i}+t_{x}-u_{1}^{i})^{2}\,,$ (A.2)
$f_{v}^{i}=(r_{21}u_{2}^{i}+r_{22}v_{2}^{i}+r_{23}z_{2}^{i}+t_{y}-v_{1}^{i})^{2}\,,$ (A.3)
and $r_{ij}$ is the $(i,j)$ element of the relative rotation $\mathrm{R_{21}}$. Please see Harris [6] for details. Here we derive a relationship between the unresolvable parameter $\eta$ and corresponding depth estimates (surface geometry) ${z_{2}^{i}}$. As the depths that minimize Eq. A.1 should satisfy
$\frac{\partial}{\partial z_{2}^{i}}f_{\mathrm{IM}}^{\prime}=0\,,$ (A.4)
@@
$z_{2}^{i}=-\frac{r_{13}(\tilde{u}^{i}+t_{x})+r_{23}(\tilde{v}^{i}+t_{y})}{r_{13}^{2}+r_{23}^{2}}\,,$ (A.5)
@@
$\tilde{u}^{i}=r_{11}u_{2}^{i}+r_{12}v_{2}^{i}-u_{1}^{i}\,,$ (A.6)
$\tilde{v}^{i}=r_{21}u_{2}^{i}+r_{22}v_{2}^{i}-v_{1}^{i}\,.$ (A.7)
Relationships between the elements of $\mathrm{R_{21}}$ and corresponding Euler angles $(\phi,\eta,\theta)$ are
$r_{11}$ $=\cos\phi\cos\theta+\sin\phi\sin\theta\cos\eta\,,$ (A.8)
$r_{12}$ $=\cos\phi\sin\theta-\sin\phi\cos\theta\cos\eta\,,$ (A.9)
$r_{13}$ $=\sin\phi\sin\eta\,,$ (A.10)
$r_{21}$ $=\sin\phi\cos\theta-\cos\phi\sin\theta\cos\eta\,,$ (A.11)
$r_{22}$ $=\sin\phi\sin\theta+\cos\phi\cos\theta\cos\eta\,,$ (A.12)
$r_{23}$ $=-\cos\phi\sin\eta\,,$ (A.13)
$r_{31}$ $=-\sin\theta\sin\eta\,,$ (A.14)
$r_{32}$ $=\cos\theta\sin\eta\,,$ (A.15)
$r_{33}$ $=\cos\eta\,.$ (A.16)
By substituting these for Eq. A.5, we have
$z_{2}^{i}=\frac{1}{\tan\eta}n_{\theta}^{i}-\frac{1}{\sin\eta}(n_{\phi}^{i}-n_{t,\phi})\,,$ (A.17)
@@
$n_{\theta}^{i}=u_{2}^{i}\sin\theta-v_{2}^{i}\cos\theta\,,$ (A.18)
$n_{\phi}^{i}=u_{1}^{i}\sin\phi-v_{1}^{i}\cos\phi\,,$ (A.19)
$n_{t,\phi}=t_{x}\sin\phi-t_{y}\cos\phi\,.$ (A.20)
@@
$\frac{1}{\sin\hat{\eta}}=\frac{\sin\eta}{\sin\hat{\eta}}\frac{1}{\sin\eta}\,,$ (A.21)
$\frac{1}{\tan\hat{\eta}}=\frac{\sin\eta}{\sin\hat{\eta}}\frac{1}{\tan\eta}+\left(\frac{\cos\hat{\eta}}{\cos\eta}-1\right)\frac{\sin\eta}{\sin\hat{\eta}}\frac{1}{\tan\eta}\,.$ (A.22)
By using Eqs. A.17, A.21 and A.22, we can relate (inaccurate) depth estimates $\hat{z}_{2}^{i}$ that correspond to $\hat{\eta}$ with the ground truth $z_{2}^{i}$
$\hat{z}_{2}^{i}=\lambda_{\hat{\eta}}z_{2}^{i}+\mu_{\hat{\eta}}u_{2}^{i}+\nu_{\hat{\eta}}v_{2}^{i}+c_{\hat{\eta}}\,,$ (A.23)
@@
$\lambda_{\hat{\eta}}=\frac{\sin\eta}{\sin\hat{\eta}}\,,$ (A.24)
$\mu_{\hat{\eta}}=\left(\frac{\cos\hat{\eta}}{\cos\eta}-1\right)\frac{\sin\eta}{\sin\hat{\eta}}\frac{1}{\tan\eta}\sin\theta\,,$ (A.25)
$\nu_{\hat{\eta}}=-\left(\frac{\cos\hat{\eta}}{\cos\eta}-1\right)\frac{\sin\eta}{\sin\hat{\eta}}\frac{1}{\tan\eta}\cos\theta\,,$ (A.26)
$c_{\hat{\eta}}=\frac{1}{\sin\hat{\eta}}\left(\hat{n}_{t,\phi}-n_{t,\phi}\right)\,,$ (A.27)
@@
$\hat{\mathrm{G}}_{2}=\mathrm{G}_{2}\mathrm{G}_{\hat{\eta},2}^{-1}\,,$ (A.29)

### v2

@@ 3.5 Camera Pose Estimation from Two Images
Based on the algorithm in Sec. 3.4, we can derive a method, as depicted in Figure 4, that takes in only two-view images of a textureless, non-Lambertian object and estimates the relative camera pose (the relative rotation and the relative translation) between the two views. We achieve this by recovering a surface normal and a reflectance map for each view, detecting the correspondences from the recovered maps, and applying the proposed algorithm to the detected correspondences.
To recover the normal and reflectance maps for each view separately, we use DeepShaRM [36] which is a radiometry-based geometry estimation method. DeepShaRM can jointly recover geometry, reflectance maps, and normal maps from an image by alternating between (1) learning-based reflectance map estimation from images and a geometry estimate, (2) learning-based surface normal estimation from images and the estimated reflectance maps, and (3) geometry optimization with the estimated surface normals. This gives us the normal and reflectance maps but only up to the GBR transformations. Please see the supplementary material for more implementation details.
To obtain 3D correspondences in the normal maps, we train a convolutional neural network to extract pixel-wise, view-invariant features that we use to match surface points between normal maps. We also train another deep neural network for the detection of reflection correspondences in the same way. We train the network by contrastive learning. We use pairs of synthetic normal maps of two views as training data. We feed both of them to the feature extraction network and obtain corresponding feature maps ${\bf F}_{1}$ and ${\bf F}_{2}$. Given ground truth correspondences $(u_{1}^{i},v_{1}^{i})\leftrightarrow(u_{2}^{i},v_{2}^{i})$, we impose the InfoNCE loss [25]
@@
The training with synthetic normal maps is, however, insufficient in practice as the surface normal maps estimated by the single-view estimation are distorted by GBR transform. We overcome this with data augmentation. For each view in the training data, we randomly sample parameters of the GBR transformation and transform the input normal map according to Eq. 4. We use the transformed normal maps instead of the original ones so that the network can learn to extract features robust to the estimation errors caused by the bas-relief ambiguity.
At inference time, using the extracted feature maps, we detect correspondences by brute-force matching and filter them with the ratio test by Lowe [20]. For any location on the object in the first image (or reflectance map), we look for the best and second best locations in the second image (or reflectance map) in terms of the cosine similarity of the features. If the ratio of the similarity of the second best location to one of the best location is lower than a threshold, we use the best location as the matched location. This gives us all of pixel, 3D, and reflection correspondences.
Once we have the three types of correspondences, we can recover the relative rotation and the relative translation using the two-step estimation algorithm (Sec. 3.4).
@@ 3.6 Joint Shape and Camera Pose Recovery
Once we obtain the relative camera pose, we can further improve its accuracy by consolidating the multi-view surface normal estimates using the camera pose estimates and exploiting the improved surface normals and corresponding reflectance maps to update the camera pose estimate. As DeepShaRM [36] is originally designed for posed multi-view images, we achieve this by alternating between multi-view surface normal and reflectance map estimation by DeepShaRM and camera pose estimation by our method. As a byproduct of DeepShaRM, we can also obtain an accurate object shape.
@@ 4 Experimental Results
Table 1: (4(a)) Mean camera pose (relative rotation) estimation accuracy on images in the nLMVS-Synth dataset [35]. (4(b)) Camera pose estimation accuracy on real-world images. (*) COLMAP [30] uses 11 view uncropped images as inputs.
@@
w/o Data Augm. 7.4 deg
w/o Joint 6.2 deg
w/o RM 6.6 deg
Ours 4.5 deg
(a) Synthetic Result
Planck Horse Bunny Cat
COLMAP [30] (*) 15.7 deg N/A 24.5 deg N/A
LightGlue [17] 5.0 deg 18.9 deg 93.6 deg 35.4 deg
SAMURAI [3] 5.6 deg 24.3 deg 30.5 deg 68.7 deg
Ours 1.9 deg 12.5 deg 7.2 deg 8.6 deg
(b) Real Data Result
We focus our experiments on answering the following key questions. How well do the reflection correspondences contribute to the camera pose estimation accuracy? Can the novel feature extraction network establish correspondences robust to GBR distortion? How does our method generalize to real-world images? How does our method compare to structure-from-motion methods and neural image synthesis methods with camera pose optimization? To answer these questions, we validate the following points
@@
  • The proposed components (especially the reflection correspondences) are essential for accurate reconstruction from a sparse set (2) of images.
  • Our method can recover accurate camera poses even from two real images.
Training Data We trained the deep networks in our method (including DeepShaRM [36]) on the training set of the nLMVS-Synth dataset [35]. The training set consists of 26850 images of 2685 synthetic objects. The training shapes are composed of primitive shapes (ellipsoids, cubes, cylinders) augmented with random height fields [34]. 94 materials and 2685 environmental maps from existing databases [21, 10, 38] are used for rendering.
Figure 5: Evaluation on synthetic shapes with different levels of flatness. (5(a)) We create such shapes by applying GBR transformations with different $\lambda$. (5(a)) Accuracy for each $\lambda$. The results shows the effectiveness of using reflection correspondences. Please see the text for details.
4.1 Evaluation on Synthetic Data
We evaluate the accuracy of the joint camera pose and object shape estimation framework on synthetic images rendered with unseen shapes, BRDFs, and illumination maps. We used synthetic images from the test set of the nLMVS-Synth dataset [35] and some rendered by ourselves for this evaluation. In total, the test set consists of 27 combinations of 3 shapes, 3 BRDFs, and 3 illumination environments.
@@ 4.1 Evaluation on Synthetic Data
We evaluate the accuracy of our method by computing the geodesic distance between the estimated relative rotation and the ground truth. Note that evaluation of the estimated translation under orthographic projection is difficult due to unsolvable offsets regarding the two viewing directions. We compare our camera pose estimation accuracy with those by its own ablated variants, “w/o Data Augm.”, “w/o Joint,” and “w/o RM.”. “w/o Data Augm.” uses the feature extraction networks trained without the data augmentation with GBR transforms. “w/o Joint” is our method without the joint iterative estimation which recovers camera poses from only the initial estimates of normal and reflectance maps. “w/o RM” ignores the bas-relief ambiguity and recovers the relative rotation using only pixel and 3D correspondences based on Eqs. 1 and 3.
@@
Figure 4(a) shows the average camera pose estimation errors. The results show that the joint estimation, the data augmentation method, and the use of reflection correspondences are essential for accurate camera pose estimation.
4.2 Robustness to Various Levels of Flatness
To further clarify the effectiveness of the reflection correspondences, we also test our method on synthetic shapes with different levels of flatness. As shown in Figure 5(a), we created such test shapes by applying GBR transformations with different parameters to the Stanford Bunny [15]. We set $\mu$ and $\nu$ in the GBR transformations to be zero and $\lambda$ to be one of 0.69, 0.83, 1.0, 1.2, and 1.44. Using the distorted shapes along with BRDF data and illumination maps used in the synthetic evaluation above, we rendered synthetic images for this evaluation. We tested our method and the baseline method that does not exploit reflection correspondences (“w/o RM”) on them.
Figure 5(b) shows the mean camera pose estimation error for each $\lambda$, i.e., each level of flatness. The baseline method that exploits only conventional correspondences works well only for a shape with a “normal” level of flatness (i.e., close to $\lambda=1$). This is because the method heavily relies on a realistic prior learned by the deep geometry estimation method [36]. In contrast, our method is robust to various levels of flatness which is critical for practical use in the real-world.
Figure 6: Reflectance maps (RMs), normal maps, surface geometry, and relative camera poses recovered from two-view real-world images [35].
Figure 7: All types of correspondences detected from the real image pair for Fig. 6.
4.3 Evaluation on Real Data
Quantitative evaluation of our method on in-the-wild images is difficult as they are by definition difficult to register with traditional methods. Nevertheless, we were able to evaluate our method on the real images from the nLMVS-Real dataset [35] and some captured and calibrated by ourselves.
We applied our method trained on the synthetic data to images of four real objects without any fine tuning. We compared our results with those of applying COLMAP [30], LightGlue [17], and SAMURAI [3]. We found that COLMAP completely fails on our inputs, i.e., two-view cropped images that cover only the target object. For images from the nLMVS-Real, we used 11 view uncropped images that capture not only the target object but also textured object around the target (i.e., ChArUco boards) as inputs to COLMAP. Note that the uncropped images are provided by the authors of the dataset [35] as RAW data.
Figure 4(b) and Fig. 6 show quantitative and qualitative results. Note that COLMAP completely failed on the “Horse” object. As SAMURAI failed to extract a 3D mesh model from their volumetric geometry representation, we only show a normal map for SAMURAI. In contrast to these existing methods which fail on these challenging inputs, our method successfully recovers plausible camera poses, surface geometry, and reflectance maps.
4.3.1 Scenes with textured and textureless objects
We can also leverage our method to estimate camera poses from sparse images capturing scenes consisting of both textured and textureless, non-Lambertian objects often found in the real world. In such a setting, the pixel correspondences can come from the textured object and the reflection correspondences from the textureless object. That is, the different types of correspondences can come from different objects in the scene to recover the unique camera poses capturing all the objects
Figure 8: Results on images capturing scenes consisting of both textured and textureless, non-Lambertian objects. Our result aligns almost perfectly with the ground truth.
in each shot.
Figure 8 shows qualitative result of our method and that of a baseline structure-from-motion (SfM) method on an example scene. Note that, although the cameras are almost orthographic, we can still use a perspective model to obtain a unique estimate of the relative rotation from only the pixel correspondences. The results show that, while the SfM method (orange) struggles to recover the angle between the two views (i.e., $\eta$), our camera pose estimation result (green) is almost identical to the ground truth (blue). This demonstrates the effectiveness of our method even in scenes consisting of multiple different types of objects. Please see the supplementary material for more results.
@@ 5 Conclusion
We introduced a new type of correspondences, reflection correspondences, which enables accurate and robust camera pose and joint shape estimation from the appearance alone of textureless, reflective objects. We believe this new type of correspondences not only expunges restricting requirements on image capture for camera pose estimation but also opens new use of object appearance in various applications. We hope to explore these in future work and also catalyze research for this with our release of code and data.
Acknowledgements
This work was in part supported by JSPS KAKENHI 20H05951 and 21H04893; JST JPMJCR20G7, JPMJAP2305, and JPMJSP2110; RIKEN GRP; ANR project TOSAI ANR-20-IADJ-0009; and the ERC grant "explorer" (No. 101097259).
Appendix 0.A Evaluation on Synthetic Data
Figure A.9 shows inlier correspondences detected by our method. Most of them join points on the surface and reflectance maps that visually appear to be in correspondence. These results demonstrate the effectiveness of the feature extractor and the robust estimation method. Their accuracy is ultimately reflected in the accuracy of the resulting camera pose estimates.
Figure A.10 shows the mean camera pose (relative rotation) estimation error for each illumination, each material, and each shape in the synthetic test data. Our results are at least comparable to and in most cases more accurate than those of other methods. These results show the effectiveness and robustness of our full method.
Figure A.9: Examples of correspondences detected from the synthetic test images. Visual inspection shows that they match semantically correct surface and reflectance map points.
(a) Accuracy for each illumination
(b) Accuracy for each material
(c) Accuracy for each shape
Figure A.10: Mean camera pose (relative rotation) estimation error for each illumination, each material, and each shape in the synthetic test data. The results show the effectiveness and the robustness of our full method. Please see Fig. I.21 for the names of the illuminations, the materials, and the shapes.
Appendix 0.B Evaluation on Real Data
Figure B.11: Qualitative results on real image pairs from the nLMVS-Real dataset [35] and on an image pair captured and calibrated by ourselves (the third object).
Figure B.12: Correspondences detected from the real image pairs in the nLMVS-Real dataset [35] and on an image pair captured by ourselves.
Figures B.11 and B.12 show qualitative results on real image pairs from the nLMVS-Real dataset [35] and one captured by ourselves. Our method successfully recovers plausible camera poses, surface geometry, surface normals, reflectance maps, and correspondences across views even for real objects without requiring textures and overlapping background.
0.B.0.1 Comparison with DUSt3R [33]
We also compare our method with DUSt3R. As the output of DUSt3R is a point map and relative camera pose estimation from it does not resolve the overall scale, we can only achieve qualitative comparison. Figure B.13 shows the results. The geometry recovered with DUSt3R is coarse. In contrast, our method successfully recovers details of the surface (e.g., bumps of the Bunny).
Figure B.13: Qualitative comparison with DUSt3R [33]. The geometry recovered with DUSt3R is coarse. In contrast, our method successfully recovers details of the surface (e.g., bumps of the Bunny).
0.B.1 Scenes with Textured and Textureless Objects
Figure B.14: Reflectance maps and camera poses recovered from two-view real images of scenes with textured and textureless, non-Lambertian objects. Leveraging different types of correspondences from different objects in the same scene leads to more accurate camera pose estimation.
Figure B.15: Normal maps and surface geometry of textureless, non-Lambertian objects recovered from the two-view real images of scenes with different objects. The recovered geometry is qualitatively plausible for the captured objects.
Figure B.16: Correspondences detected from the real images of scenes with textured and textureless, non-Lambertian objects.
Figures B.14, B.15 and B.16 show camera poses, reflectance maps, normal maps, surface geometry, and correspondences estimated from two-view real images of scenes with textured and textureless, non-Lambertian objects. By leveraging different types of correspondences from different objects in the same scene, we can recover accurate camera poses and plausible surface geometry for complex in-the-wild scenes.
Appendix 0.C Ambiguity in The Estimation with Conventional Correspondences
In this section, we prove that, under orthographic projection (e.g., cameras are distant from the object), relative camera pose estimation between two views suffers from a fundamental ambiguity that cannot be resolved even when correct pixel and 3D correspondences are provided.
@@ Appendix 0.C Ambiguity in The Estimation with Conventional Correspondences
$\min_{z_{2}^{i},t_{x},t_{y}}f_{\mathcal{I}}^{\prime}=\min_{z_{2}^{i},t_{x},t_{y}}\sum_{i}f_{u}^{i}+f_{v}^{i}\,,$ (0.C.1)
where $z_{2}^{i}$ is the depth of a surface point in the second view for the $i$-th correspondence, $t_{x}$ and $t_{y}$ are the $x$ and $y$ elements of the relative translation vector, respectively, $f_{u}^{i}$ and $f_{v}^{i}$ are reprojection errors along the x and y axes, respectively,
$f_{u}^{i}=(r_{11}u_{2}^{i}+r_{12}v_{2}^{i}+r_{13}z_{2}^{i}+t_{x}-u_{1}^{i})^{2}\,,$ (0.C.2)
$f_{v}^{i}=(r_{21}u_{2}^{i}+r_{22}v_{2}^{i}+r_{23}z_{2}^{i}+t_{y}-v_{1}^{i})^{2}\,,$ (0.C.3)
and $r_{ij}$ is the $(i,j)$ element of the relative rotation $\mathrm{R_{21}}$. Please see Harris [12] for details.
Let us derive the relationship between the unresolvable parameter $\eta$ and corresponding depth estimates (surface geometry) ${z_{2}^{i}}$. As the depth values that minimize Eq. 0.C.1 should satisfy
$\frac{\partial}{\partial z_{2}^{i}}f_{\mathcal{I}}^{\prime}=0\,,$ (0.C.4)
@@
$z_{2}^{i}=-\frac{r_{13}(\tilde{u}^{i}+t_{x})+r_{23}(\tilde{v}^{i}+t_{y})}{r_{13}^{2}+r_{23}^{2}}\,,$ (0.C.5)
@@
$\tilde{u}^{i}=r_{11}u_{2}^{i}+r_{12}v_{2}^{i}-u_{1}^{i}\,,$ (0.C.6)
$\tilde{v}^{i}=r_{21}u_{2}^{i}+r_{22}v_{2}^{i}-v_{1}^{i}\,.$ (0.C.7)
The elements of $\mathrm{R_{21}}$ are related to their corresponding Euler angles $(\phi,\eta,\theta)$ by
$r_{11}$ $=\cos\phi\cos\theta+\sin\phi\sin\theta\cos\eta\,,$ (0.C.8)
$r_{12}$ $=\cos\phi\sin\theta-\sin\phi\cos\theta\cos\eta\,,$ (0.C.9)
$r_{13}$ $=\sin\phi\sin\eta\,,$ (0.C.10)
$r_{21}$ $=\sin\phi\cos\theta-\cos\phi\sin\theta\cos\eta\,,$ (0.C.11)
$r_{22}$ $=\sin\phi\sin\theta+\cos\phi\cos\theta\cos\eta\,,$ (0.C.12)
$r_{23}$ $=-\cos\phi\sin\eta\,,$ (0.C.13)
$r_{31}$ $=-\sin\theta\sin\eta\,,$ (0.C.14)
$r_{32}$ $=\cos\theta\sin\eta\,,$ (0.C.15)
$r_{33}$ $=\cos\eta\,.$ (0.C.16)
By substituting these for Eq. 0.C.5, we have
$z_{2}^{i}=\frac{1}{\tan\eta}n_{\theta}^{i}-\frac{1}{\sin\eta}(n_{\phi}^{i}-n_{t,\phi})\,,$ (0.C.17)
@@
$n_{\theta}^{i}=u_{2}^{i}\sin\theta-v_{2}^{i}\cos\theta\,,$ (0.C.18)
$n_{\phi}^{i}=u_{1}^{i}\sin\phi-v_{1}^{i}\cos\phi\,,$ (0.C.19)
$n_{t,\phi}=t_{x}\sin\phi-t_{y}\cos\phi\,.$ (0.C.20)
@@
$\frac{1}{\sin\hat{\eta}}=\frac{\sin\eta}{\sin\hat{\eta}}\frac{1}{\sin\eta}\,,$ (0.C.21)
$\frac{1}{\tan\hat{\eta}}=\frac{\sin\eta}{\sin\hat{\eta}}\frac{1}{\tan\eta}+\left(\frac{\cos\hat{\eta}}{\cos\eta}-1\right)\frac{\sin\eta}{\sin\hat{\eta}}\frac{1}{\tan\eta}\,.$ (0.C.22)
By using Eqs. 0.C.17, 0.C.21 and 0.C.22, we can relate (inaccurate) depth estimates $\hat{z}_{2}^{i}$ that correspond to $\hat{\eta}$ with the ground truth $z_{2}^{i}$
$\hat{z}_{2}^{i}=\lambda_{\hat{\eta}}z_{2}^{i}+\mu_{\hat{\eta}}u_{2}^{i}+\nu_{\hat{\eta}}v_{2}^{i}+c_{\hat{\eta}}\,,$ (0.C.23)
@@
$\lambda_{\hat{\eta}}=\frac{\sin\eta}{\sin\hat{\eta}}\,,$ (0.C.24)
$\mu_{\hat{\eta}}=\left(\frac{\cos\hat{\eta}}{\cos\eta}-1\right)\frac{\sin\eta}{\sin\hat{\eta}}\frac{1}{\tan\eta}\sin\theta\,,$ (0.C.25)
$\nu_{\hat{\eta}}=-\left(\frac{\cos\hat{\eta}}{\cos\eta}-1\right)\frac{\sin\eta}{\sin\hat{\eta}}\frac{1}{\tan\eta}\cos\theta\,,$ (0.C.26)
$c_{\hat{\eta}}=\frac{1}{\sin\hat{\eta}}\left(\hat{n}_{t,\phi}-n_{t,\phi}\right)\,,$ (0.C.27)
@@
$\hat{\mathrm{G}}_{2}=\mathrm{G}_{2}\mathrm{G}_{\hat{\eta},2}^{-1}\,,$ (0.C.29)